a wheel makes 3 9/10 revolution per minute. if it rotates for 38 minutes how many revolutions does it make

Answer by jim_thompson5910(35256)

3 9/10 = 3 + 9/10
3 9/10 = 30/10 + 9/10
3 9/10 = 39/10
-------------------------------------------------------

The wheel makes 39/10 of a revolution per minute

(39 rotations/10 minutes)*(38 minutes) = (39/10)*38 = 148.2 revolutions